OpenStudy (anonymous):

how do i put 60.0625 in fraction form

OpenStudy (anonymous):

Daisyboo23 do you have a graphing calculator?

OpenStudy (texaschic101):

60 625/10,000

OpenStudy (anonymous):

there are four digits after decimal so write 1 and four digits in the denominator. \[\frac{ 600625 }{ 10,000 }\]

OpenStudy (texaschic101):

I just made mine into a mixed number....it can be turned into an improper fraction

OpenStudy (unklerhaukus):

0.0625=1/16

OpenStudy (anonymous):



OpenStudy (anonymous):

\[ \frac{600625}{10000}=\frac{961}{16} \]

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=60.0625